# HugAgent: Evaluating LLMs in Simulating Human-Like Individual Reasoning on Open-Ended Tasks

### v1

@@ 3.4 Dataset Statistics
We additionally generate synthetic agents with scripted causal belief graphs and deterministic update operators. In total, the synthetic track comprises $50$ agents with a total of 1,4781{,}478 items, providing controlled coverage and stress tests. Details of synthetic construction are given in Appendix N.
Human Track (N=54) Synthetic Track (N=50)
@@
Belief-State Inf. 108 122 126 356 79 92 122 293
Belief Dyn. Update 472 364 550 1,386 458 324 403 1,185
Total Items 580 486 676 1,742 537 416 525 1,478
Table 1: Dataset statistics by task and domain. Human track (N=54 participants) provides ecological validity; synthetic track (N=50 scripted agents) scales coverage.
@@ 3.5 Evaluation Protocols
We evaluate belief state inference using accuracy, the proportion of exact matches with ground-truth labels. For belief dynamic update, we report three metrics: (i) accuracy, the proportion of predictions within a tolerance band of the true response ($\pm$1 for 5-point, $\pm$2 for 10-point scales); (ii) mean absolute error (MAE), the average magnitude of deviation from the ground truth, with all responses normalized to a 5-point scale for consistency; and (iii) directional accuracy, the proportion of updates where the predicted change matched the true change in sign (increase, decrease, or no change). Formal definitions are provided in Appendix Q.1.
@@
We re-contacted a subset of participants for a short-interval (14-day) test–retest study. Across all sessions, 54 participants contributed data. Of these, 18 completed the retest, and 10 were retained following a demographic consistency check. Belief State Inference yielded an accuracy of 83.10% (SD = 0.37, 95% CI: [78.7, 87.5]). Belief Dynamic Update achieved an accuracy of 88.22% (SD = 1.21, 95% CI: [85.4, 90.6]) and a mean absolute error of 0.62 (SD = 0.79, 95% CI: [0.56, 0.69]). These scores establish a human consistency ceiling, as outlined in Section 3.1, against which model performance can be benchmarked.
@@ 6 Why it Happens: Ablations, Error Analysis and Mitigation Strategies
Task Method Accuracy (% $\uparrow$) MAE ($\downarrow$) Healthcare Surveillance Zoning Avg Healthcare Surveillance Zoning Avg Belief-State Inference GPT-4o (No-Context) 55.93±0.01 66.07±0.01 53.49±0.02 58.50±0.02 N/A GPT-4o (Full) 71.11±0.01 75.25±0.01 77.62±0.00 74.66±0.01 N/A Qwen2.5-32B (No-Context) 45.19±0.02 59.02±0.01 48.57±0.01 50.93±0.01 N/A Qwen2.5-32B (Full) 77.96±0.00 76.56±0.00 76.98±0.01 77.17±0.01 N/A Belief Dynamic Update GPT-4o (No-Context) 30.89±0.08 48.21±0.00 36.45±0.00 39.77±0.06 1.77±0.01 1.51±0.01 1.84±0.01 1.71±0.01 GPT-4o (Full) 61.57±0.00 63.74±0.01 64.04±0.00 63.12±0.00 1.27±0.01 1.35±0.01 1.26±0.01 1.29±0.01 Qwen2.5-32B (No-Context) 34.58±0.01 33.90±0.01 27.89±0.00 32.12±0.01 1.89±0.01 1.75±0.01 1.99±0.01 1.88±0.01 Qwen2.5-32B (Full) 56.82±0.00 57.80±0.00 62.25±0.00 58.96±0.00 1.43±0.00 1.43±0.01 1.33±0.01 1.40±0.01
@@
Method Belief-State Inference Belief Dynamic Update Accuracy (% $\uparrow$) Accuracy (% $\uparrow$) MAE ($\downarrow$) MSE ($\downarrow$) GPT-4o 54.44±0.06 39.30±0.03 1.93±0.10 N/A Qwen2.5-32B 60.44±0.03 37.61±0.04 2.03±0.14 N/A
@@ 6.3 Error Analysis: Bias Patterns and Sources of Failure
We observe systematic, topic-conditioned biases beyond metrics.
@@
Performance differs by topic. Within-domain, belief-state inference peaks in Surveillance for both models (LLaMA 3.3 70B 80.66%; GPT-5-mini 76.89%). For update, peaks differ by model: LLaMA 3.3 70B is best on Zoning (73.09%, MAE 1.14), whereas GPT-5-mini is best on Healthcare (59.92%, MAE 1.37).
Cross-topic transfer degrades performance. Using Surveillance context to predict Zoning, belief dynamic updates reduces Qwen2.5-32B-instr from its in-domain Zoning result (62.25%; MAE 1.42) to 24.84% (MAE 1.94). For GPT-4o, the in-domain Zoning update is 64.04% (MAE 1.26); with Surveillance context it drops to 42.47% (MAE 1.60).
@@ 7 Mitigation Strategies: Insights and Guiding Principles
Our error analysis revealed structured failure modes that recur across model families. Rather than ad-hoc fixes, we position these as diagnostic insights motivating guiding principles for future work: (i) Topic sensitivity. Errors vary sharply across healthcare, zoning, and surveillance. Principle: apply calibration (offset correction, domain-specific regularization) to align predictions with human heterogeneity.
These principles underscore HugAgent’s role not only as a benchmark but also as a diagnostic tool, surfacing structured biases that standard metrics (accuracy, MAE) obscure. Appendix R provides proof-of-concept demonstrations. The larger contribution, however, is to chart a research agenda: HugAgent offers a reproducible setting in which such mitigation strategies can be systematically developed and evaluated.
@@ Appendix C HugAgent Benchmark Details
Belief State Inference yielded an accuracy of 83.10% (SD = 0.37, 95% CI: [78.7, 87.5]). By topic, accuracies were 85.56% for Surveillance (SD = 0.35, 95% CI: [78.3, 92.8]), 82.11% for Zoning (SD = 0.38, 95% CI: [74.4, 89.8]), and 81.82% for Healthcare (SD = 0.39, 95% CI: [74.2, 89.4]).
Belief Dynamic Update achieved an accuracy of 88.22% (SD = 1.21, 95% CI: [85.4, 90.6]) and a mean absolute error of 0.62 (SD = 0.79, 95% CI: [0.56, 0.69]). Across topics, accuracy was 87.08% for Healthcare (SD = 1.26), 91.67% for Surveillance (SD = 1.09), and 86.21% for Zoning (SD = 1.25). Corresponding mean absolute errors were 0.65 (SD = 0.82, 95% CI: [0.55, 0.76]) for Healthcare, 0.56 (SD = 0.62, 95% CI: [0.47, 0.65]) for Surveillance, and 0.64 (SD = 0.90, 95% CI: [0.50, 0.77]) for Zoning.
@@
Model Human(N=54) Synthetic(N=50)
@@
Health Surveil. Zoning Health Surveil. Zoning Health Surveil. Zoning Health Surveil. Zoning
@@
GPT-4o 71.11±0.77 75.25±0.90 77.62±0.35 61.57±0.19, 1.27±0.01 63.74±0.58, 1.35±0.01 64.04±0.20, 1.26±0.01 60.76±0.00 75.00±0.77 68.03±0.58 52.71±0.29, 1.72±0.00 65.00±0.35, 1.18±0.01 58.11±0.48, 1.41±0.01
GPT-5-mini 73.15±2.36 76.89±0.90 75.87±0.43 59.92±1.05, 1.37±0.02 55.16±1.09, 1.54±0.02 59.56±0.73, 1.40±0.01 50.38±1.88 72.39±1.24 53.61±26.80 51.09±0.91, 1.74±0.01 64.44±0.46, 1.21±0.02 56.23±1.63, 1.46±0.02
o3-mini 69.26±1.21 77.38±1.70 78.73±0.87 60.08±0.28, 1.31±0.02 62.53±1.16, 1.29±0.03 71.02±0.79, 1.07±0.01 58.23±1.55 78.48±1.19 72.30±1.47 46.90±0.81, 1.80±0.02 58.46±0.28, 1.29±0.00 64.07±0.37, 1.26±0.01
Gemini 2.0 Flash 63.89±0.00 68.20±0.37 77.78±0.00 56.23±0.12, 1.39±0.00 62.25±0.37, 1.33±0.00 63.16±0.10, 1.33±0.00 51.90±0.00 63.48±1.24 63.61±0.45 45.28±0.19, 1.77±0.00 62.96±0.00, 1.18±0.00 55.93±0.22, 1.49±0.00
LLaMA_3.3_70B 70.74±0.51 80.66±0.45 77.78±0.00 64.28±0.38, 1.31±0.01 65.33±0.23, 1.29±0.01 73.09±0.31, 1.14±0.01 61.27±1.13 76.09±0.00 77.87±0.00 51.44±0.65, 1.82±0.01 57.65±0.34, 1.32±0.00 65.96±0.21, 1.32±0.00
DeepSeek-R1-0528 76.11±0.41 75.41±1.16 74.76±0.66 61.44±0.58, 1.33±0.00 68.30±0.63, 1.27±0.02 64.91±0.41, 1.26±0.01 62.78±1.13 75.87±1.42 73.61±0.69 60.52±0.18, 1.59±0.01 64.75±0.40, 1.15±0.01 58.21±0.45, 1.40±0.01
Qwen-max_2024-10-15 78.52±0.77 76.39±0.37 77.30±0.43 56.78±0.64, 1.43±0.01 57.69±0.27, 1.44±0.01 62.11±0.35, 1.34±0.00 62.78±2.12 68.48±0.00 71.97±0.37 45.20±0.44, 1.80±0.00 64.57±0.14, 1.18±0.01 55.83±0.46, 1.50±0.01
Qwen-plus_2025-07-28 78.70±0.65 76.39±1.22 77.62±0.66 56.74±0.59, 1.44±0.01 57.86±0.15, 1.43±0.01 62.18±0.18, 1.34±0.00 63.80±1.44 67.83±1.97 71.80±0.73 45.11±0.37, 1.80±0.00 64.94±0.17, 1.18±0.01 55.83±0.53, 1.50±0.01
Qwen2.5-32B-instr 77.96±0.41 76.56±0.45 76.98±0.79 56.82±0.41, 1.43±0.00 57.80±0.25, 1.43±0.01 62.25±0.16, 1.33±0.01 64.30±0.56 68.70±0.49 71.64±0.93 45.24±0.39, 1.80±0.00 65.00±0.17, 1.18±0.00 55.88±0.54, 1.50±0.01
Qwen2.5-7B-instr 78.33±0.51 75.74±2.06 77.46±0.43 56.82±0.41, 1.43±0.01 57.64±0.45, 1.43±0.01 62.00±0.34, 1.34±0.01 63.80±1.13 68.48±1.33 71.15±0.90 45.11±0.29, 1.80±0.00 64.81±0.22, 1.18±0.01 56.08±0.25, 1.50±0.01
Generative Agents (Park et al., 2023) 76.11±0.41 77.21±0.37 75.24±0.66 57.67±0.76, 1.43±0.01 55.66±0.69, 1.44±0.01 61.35±0.38, 1.34±0.00 68.61±0.57 77.39±0.91 74.43±0.37 46.24±0.68, 1.79±0.01 64.26±0.26, 1.17±0.00 55.78±0.21, 1.52±0.01
RAG (Lewis et al., 2020) 71.67±0.51 78.36±1.37 76.35±0.35 50.64±0.58, 1.60±0.01 53.68±0.60, 1.50±0.00 51.38±0.24, 1.60±0.00 51.39±1.44 67.61±0.90 70.16±1.37 43.49±0.28, 1.85±0.00 61.98±0.40, 1.22±0.01 50.07±0.32, 1.61±0.00
RAG-FC (Lewis et al., 2020) 77.22±0.83 76.72±0.93 78.73±0.87 58.73±0.55, 1.43±0.01 59.34±0.34, 1.41±0.01 61.85±0.15, 1.33±0.01 68.10±1.65 76.09±1.09 74.26±1.10 53.41±0.45, 1.70±0.01 64.26±0.46, 1.20±0.01 60.35±0.59, 1.44±0.01
@@
Table 8: Results on Human and Synthetic datasets for belief-state inference and belief dynamic update tasks. Values are mean±std over 5 runs; for belief dynamic update, both Accuracy and MAE are reported.

### v2

# HugAgent: Evaluating LLMs in Simulating Individual-Level Human Reasoning on Open-Ended Tasks

@@ 3.4 Dataset Statistics
We construct a parallel synthetic track by assigning each agent a scripted causal belief network (CBN) that defines its reasoning structure and deterministic update rules. All synthetic agents follow the same survey and interview protocol as in the human track, allowing direct comparison under identical tasks. In total, the benchmark includes 500 synthetic agents, from which we use a stratified subset of 50 agents for computation and analysis. Unless otherwise specified, all experiments reported in this paper are conducted on this subset. Both the full and subset data are included in the benchmark release. Detailed construction procedures are provided in Appendix N.
Human Track (N=54) Synthetic Track (N=500)
@@
Belief-State Inf. 108 122 126 356 1,303 1,297 1,302 3,902
Belief Dyn. Update 472 364 550 1,386 3,888 2,610 3,818 10,316
Total Items 580 486 676 1,742 5,191 3,907 5,120 14,218
Table 1: Dataset statistics by task and domain. Human track (N=54 participants) provides ecological validity; synthetic track (N=500 scripted agents) scales coverage.
@@ 3.5 Evaluation Protocols
We evaluate belief state inference using accuracy, the proportion of exact matches with ground-truth labels. For belief dynamic update, we report four metrics: (i) accuracy, the proportion of predictions within a tolerance band of the true response ($\pm$1 for 5-point, $\pm$2 for 10-point scales); (ii) mean absolute error (MAE), the average magnitude of deviation from the ground truth, with all responses normalized to a 5-point scale for consistency; (iii) directional accuracy measures whether the predicted belief update matches the ground truth in direction(increase, decrease, or no change); and (iv) unified aggregate score(UAS), following the GLUE (Wang et al., 2018) and SuperGLUE (Wang et al., 2019) scoring paradigm, we derive an overall score by averaging task-specific metrics (accuracy, MAE, directional accuracy) and then taking the unweighted mean across belief-state inference and belief-dynamics update, ensuring balanced comparability across tasks.Formal definitions and computation details are given in Appendix Q.1.
@@
We re-contacted a subset of participants for a short-interval (14-day) test–retest study. Across all sessions, 54 participants contributed data. Of these, 18 completed the retest, and 13 were retained following a demographic consistency check. Belief State Inference yielded an accuracy of 84.49% (SD = 0.36, 95% CI: [80.82, 88.16]). Belief Dynamic Update achieved an accuracy of 85.66% (SD = 1.27, 95% CI: [83.34, 87.70]) and a mean absolute error of 0.67 (SD = 0.83, 95% CI: [0.63, 0.73]). These scores establish a human consistency ceiling, as outlined in Section 3.1, against which model performance can be benchmarked.
@@ 6 Why it Happens: Ablations, Error Analysis and Mitigation Strategies
Task Method Accuracy (% $\uparrow$) MAE ($\downarrow$) Healthcare Surveillance Zoning Avg Healthcare Surveillance Zoning Avg Belief-State Inference GPT-4o (No-Context) 55.93±0.01 66.07±0.01 53.49±0.02 58.50±0.02 N/A GPT-4o (Full) 71.11±0.01 75.25±0.01 77.62±0.00 74.66±0.01 N/A Qwen2.5-32B (No-Context) 45.19±0.02 59.02±0.01 48.57±0.01 50.93±0.01 N/A Qwen2.5-32B (Full) 77.96±0.00 76.56±0.00 76.98±0.01 77.17±0.01 N/A Belief Dynamic Update GPT-4o (No-Context) 34.62±0.00 48.21±0.00 36.45±0.00 39.77±0.06 1.77±0.01 1.51±0.01 1.84±0.01 1.71±0.01 GPT-4o (Full) 61.57±0.00 63.74±0.01 64.04±0.00 63.12±0.00 1.27±0.01 1.35±0.01 1.26±0.01 1.29±0.01 Qwen2.5-32B (No-Context) 34.58±0.01 33.90±0.01 27.89±0.00 32.12±0.01 1.89±0.01 1.75±0.01 1.99±0.01 1.88±0.01 Qwen2.5-32B (Full) 56.82±0.00 57.80±0.00 62.25±0.00 58.96±0.00 1.43±0.00 1.43±0.01 1.33±0.01 1.40±0.01
@@
Method Belief-State Inference Belief Dynamic Update Accuracy (% $\uparrow$) Accuracy (% $\uparrow$) MAE ($\downarrow$) MSE ($\downarrow$) GPT-4o 54.44±0.06 39.30±0.03 1.93±0.10 N/A Qwen2.5-32B 60.44±0.03 38.16±0.03 2.02±0.13 N/A
@@ 6.3 Error Analysis: Bias Patterns and Sources of Failure
Performance varies systematically across topics. Within-domain, belief-state inference peaks in Surveillance for both models (LLaMA 3.3 70B 80.66%; GPT-5-mini 76.89%). For update, LLaMA 3.3 70B performs best on Zoning (73.09%, MAE 1.14), whereas GPT-5-mini peaks on Healthcare (59.92%, MAE 1.37).
Cross-topic transfer sharply degrades performance. Using Surveillance context to predict Zoning, belief dynamic updates for Qwen2.5-32B-instr drop from in-domain Zoning (62.25%; MAE 1.42) to 24.84% (MAE 1.94); GPT-4o falls from 64.04% (MAE 1.26) to 42.47% (MAE 1.60).
@@
Directional error decomposition.
We decompose update errors into two types: (1) wrong direction—the model predicts a belief change in the opposite direction of the human’s, and (2) change mismatch—the model fails to register that a change occurred at all. (Detailed definitions in Appendix Q.1.)
Across systems, change-mismatch errors dominate. In Healthcare, GPT-4o shows 119 mismatches vs. 5 wrong directions (DA 0.47); Qwen2.5-32B-instr records 157 vs. 5 (DA 0.31). This asymmetry holds across topics, yielding high wrong-direction accuracy ($\approx$0.93–0.98) but only moderate overall directional accuracy (0.30–0.56).
The dominance of mismatch errors indicates models preserve the sign of stance change once detected, but are insensitive to when updates should occur. Rather than inverting polarity, they often remain static despite contextual cues—suggesting weak temporal calibration or over-regularization toward “no change.” Consequently, even strong models (LLaMA 3.3 70B, GPT-4o) achieve moderate update accuracy ($\approx$60–65%) yet lag humans in Dir. Acc. by 15–20 points (Table 2). Moreover, low MAE (e.g., o3-mini, MAE 1.22) does not necessarily imply higher directional accuracy, confirming that magnitude alignment alone does not ensure correct update directionality.
Implication. LLMs are directionally stable but dynamically insensitive. Improving belief-update behavior thus requires mechanisms that better detect subtle stance shifts—context-aware memory, calibrated “no-change” thresholds, or explicit reasoning-weight tracking—rather than further tuning polarity prediction.
@@ 7 Mitigation Strategies: Insights and Guiding Principles
Our error analysis revealed structured failure modes that recur across model families. Rather than ad-hoc fixes, we position these as diagnostic insights motivating guiding principles for future work: (i) Topic sensitivity. Errors vary sharply across healthcare, zoning, and surveillance. Principle: apply calibration (offset correction, domain-specific regularization) to align predictions with human heterogeneity. (ii) Directional sensitivity. Models often preserve stance polarity but miss when updates should occur, leading to low directional accuracy. Principle: introduce temporal or contextual gating—e.g., confidence-weighted “update triggers” or explicit change-detection modules—to better capture belief shifts without overreacting to noise.
These principles underscore HugAgent’s role not only as a benchmark but also as a diagnostic tool, surfacing structured biases that standard metrics (accuracy, MAE) obscure. The larger contribution, however, is to chart a research agenda: HugAgent offers a reproducible setting in which such mitigation strategies can be systematically developed and evaluated.
@@ Appendix C HugAgent Benchmark Details
Belief State Inference yielded an accuracy of 84.49% (SD = 0.36, 95% CI: [80.82, 88.16]). By topic, accuracies were 87.50% for Surveillance (SD = 0.33, 95% CI: [81.58, 93.42]), 81.61% for Zoning (SD = 0.38, 95% CI: [74.81, 88.39]), and 84.50% for Healthcare (SD = 0.36, 95% CI: [78.25, 90.74]).
Belief Dynamic Update achieved an accuracy of 85.66% (SD = 1.27, 95% CI: [83.34, 87.70]) and a mean absolute error of 0.67 (SD = 0.83, 95% CI: [0.63, 0.73]). Across topics, accuracy was 85.75% for Healthcare (SD = 1.28), 85.33% for Surveillance (SD = 1.27), and 85.86% for Zoning (SD = 1.24). Corresponding mean absolute errors were 0.66 (SD = 0.85, 95% CI: [0.57, 0.74]) for Healthcare, 0.72 (SD = 0.74, 95% CI: [0.63, 0.80]) for Surveillance, and 0.67 (SD = 0.90, 95% CI: [0.56, 0.77]) for Zoning.
@@
Model Human Dataset (N=54)
Belief-State Inference Belief Dynamic Update
@@
Health Surveil. Zoning Health Surveil. Zoning Health Surveil. Zoning
@@
OpenAI Models
GPT-4o 71.11±0.77 75.25±0.90 77.62±0.35 61.57±0.19 63.74±0.58 64.04±0.20 1.27±0.01 1.35±0.01 1.26±0.01
GPT-5-mini 73.15±2.36 76.89±0.90 75.87±0.43 59.92±1.05 55.16±1.09 59.56±0.73 1.37±0.02 1.54±0.02 1.40±0.01
o3-mini 69.26±1.21 77.38±1.70 78.73±0.87 60.08±0.28 62.53±1.16 71.02±0.79 1.31±0.02 1.29±0.03 1.07±0.01
Other Closed-Source Models
Gemini 2.0 Flash 63.89±0.00 68.20±0.37 77.78±0.00 56.23±0.12 62.25±0.37 63.16±0.10 1.39±0.00 1.33±0.00 1.33±0.00
DeepSeek-R1-0528 76.11±0.41 75.41±1.16 74.76±0.66 61.44±0.58 68.30±0.63 64.91±0.41 1.33±0.00 1.27±0.02 1.26±0.01
Qwen-plus_2025-07-28 78.70±0.65 76.39±1.22 77.62±0.66 56.74±0.59 57.86±0.15 62.18±0.18 1.44±0.01 1.43±0.01 1.34±0.00
Qwen-max_2024-10-15 78.52±0.77 76.39±0.37 77.30±0.43 56.78±0.64 57.69±0.27 62.11±0.35 1.43±0.01 1.44±0.01 1.34±0.00
Open-Source Models
LLaMA_3.3_70B 70.74±0.51 80.66±0.45 77.78±0.00 64.28±0.38 65.33±0.23 73.09±0.31 1.31±0.01 1.29±0.01 1.14±0.01
Qwen2.5-32B-instr 77.96±0.41 76.56±0.45 76.98±0.79 56.82±0.41 57.80±0.25 62.25±0.16 1.43±0.00 1.43±0.01 1.33±0.01
Qwen2.5-7B-instr 78.33±0.51 75.74±2.06 77.46±0.43 56.82±0.41 57.64±0.45 62.00±0.34 1.43±0.01 1.43±0.01 1.34±0.01
Memory-Augmented Baselines
RAG [^1] RAG (Lewis et al., 2020), RAG-FC = RAG with Full Context (Lewis et al., 2020). 71.67±0.51 78.36±1.37 76.35±0.35 50.64±0.58 53.68±0.60 51.38±0.24 1.60±0.01 1.50±0.00 1.60±0.00
RAG-FC [^1] RAG (Lewis et al., 2020), RAG-FC = RAG with Full Context (Lewis et al., 2020). 77.22±0.83 76.72±0.93 78.73±0.87 58.73±0.55 59.34±0.34 61.85±0.15 1.43±0.01 1.41±0.01 1.33±0.01
Generative Agents [^2] Generative Agents (Park et al., 2023). 76.11±0.41 77.21±0.37 75.24±0.66 57.67±0.76 55.66±0.69 61.35±0.38 1.43±0.01 1.44±0.01 1.34±0.00
@@
Table 8: Results on Human dataset for belief-state inference and belief dynamic update tasks across three policy topics (Health, Surveillance, Zoning). Values are mean±std over 5 runs. For belief dynamic update, both Accuracy and MAE are reported separately. Best non-human results per column are in bold.
Model Synthetic Dataset (N=50)
Belief-State Inference Belief Dynamic Update
[…]
Health Surveil. Zoning Health Surveil. Zoning Health Surveil. Zoning
OpenAI Models
GPT-4o 60.76±0.00 75.00±0.77 68.03±0.58 52.71±0.29 65.00±0.35 58.11±0.48 1.72±0.00 1.18±0.01 1.41±0.01
GPT-5-mini 50.38±1.88 72.39±1.24 53.61±26.80 51.09±0.91 64.44±0.46 56.23±1.63 1.74±0.01 1.21±0.02 1.46±0.02
o3-mini 58.23±1.55 78.48±1.19 72.30±1.47 46.90±0.81 58.46±0.28 64.07±0.37 1.80±0.02 1.29±0.00 1.26±0.01
Other Closed-Source Models
Gemini 2.0 Flash 51.90±0.00 63.48±1.24 63.61±0.45 45.28±0.19 62.96±0.00 55.93±0.22 1.77±0.00 1.18±0.00 1.49±0.00
DeepSeek-R1-0528 62.78±1.13 75.87±1.42 73.61±0.69 60.52±0.18 64.75±0.40 58.21±0.45 1.59±0.01 1.15±0.01 1.40±0.01
Qwen-plus_2025-07-28 63.80±1.44 67.83±1.97 71.80±0.73 45.11±0.37 64.94±0.17 55.83±0.53 1.80±0.00 1.18±0.01 1.50±0.01
Qwen-max_2024-10-15 62.78±2.12 68.48±0.00 71.97±0.37 45.20±0.44 64.57±0.14 55.83±0.46 1.80±0.00 1.18±0.01 1.50±0.01
Open-Source Models
LLaMA_3.3_70B 61.27±1.13 76.09±0.00 77.87±0.00 51.44±0.65 57.65±0.34 65.96±0.21 1.82±0.01 1.32±0.00 1.32±0.00
Qwen2.5-32B-instr 64.30±0.56 68.70±0.49 71.64±0.93 45.24±0.39 65.00±0.17 55.88±0.54 1.80±0.00 1.18±0.00 1.50±0.01
Qwen2.5-7B-instr 63.80±1.13 68.48±1.33 71.15±0.90 45.11±0.29 64.81±0.22 56.08±0.25 1.80±0.00 1.18±0.01 1.50±0.01
Memory-Augmented Baselines
RAG11footnotemark: 1 51.39±1.44 67.61±0.90 70.16±1.37 43.49±0.28 61.98±0.40 50.07±0.32 1.85±0.00 1.22±0.01 1.61±0.00
RAG-FC11footnotemark: 1 68.10±1.65 76.09±1.09 74.26±1.10 53.41±0.45 64.26±0.46 60.35±0.59 1.70±0.01 1.20±0.01 1.44±0.01
Generative Agents22footnotemark: 2 68.61±0.57 77.39±0.91 74.43±0.37 46.24±0.68 64.26±0.26 55.78±0.21 1.79±0.01 1.17±0.00 1.52±0.01
Table 9: Results on Synthetic dataset for belief-state inference and belief dynamic update tasks across three policy topics (Health, Surveillance, Zoning). Values are mean±std over 5 runs. For belief dynamic update, both Accuracy and MAE are reported separately. Best results per column are in bold.